OpenStudy (anonymous):

ind the extreme values of the function and where they occur y = x^3 - 3x^2 + 4x - 4

zepdrix (zepdrix):

Calculus? :) So uh, differentiate, yes? Power rule? Or do you need to be using the formal definition for derivative at this point?

zepdrix (zepdrix):

Where we getting stuck?

OpenStudy (anonymous):

Yes ugh I know i have to find the derivative, set equal to zero and find the x and y But im stuck at finding the derivative Im not sure if i can use imaginary numbers in calculus

OpenStudy (anonymous):

power rule

OpenStudy (anonymous):

How would i use the power rule?

OpenStudy (anonymous):

x^3---->3x^(3-1)

OpenStudy (trojanpoem):

y' = 3x^2 - 6x + 4 b^2 - 4 ac < 0 @magepker728 , the equation has no real roots

OpenStudy (anonymous):

Graphing is the easy way out :( I need to learn how to do it algebraically

zepdrix (zepdrix):

Not too familiar with power rule just yet? You bring the exponent down in front as a multiplier, and decrease the exponent by 1. \(\large\rm \frac{d}{dx}x^n=nx^{n-1}\) So like, for your first term, \(\large\rm \frac{d}{dx}x^3=3x^{3-1}\) k? :o

OpenStudy (anonymous):

Oh, i mistook power rule for the product rule. I know how to do power rule but i guess my real problem is that i do not know how to find the x value(s) after i find the derivative

OpenStudy (trojanpoem):

@zepdrix , this is the derivative : y' = 3x^2 - 6x + 4 to find the extrema 3x^2 -6x + 4 = 0 b^2 - 4ac < 0 No real roots only complex ( a + bi ) He wanna know how to solve it while the roots aren't real., got his problem ?

OpenStudy (anonymous):

@ zepdrix... 3x^2 - 6x + 4.... I tried to factor but it didnt work then i tried the quadratic formula and i got imaginary numbers

OpenStudy (anonymous):

@trojanpoem how do i solve it if the roots aren' t? And are imaginary numbers allowed in calulus ab?

zepdrix (zepdrix):

Ok so no critical/extreme points :) cool, seems like you figured it out. Are inflection points considered extreme points? I can't remember how they categorize this stuff... Have you gone over second derivative stuff yet?

OpenStudy (trojanpoem):

@1041904 , No, Complex roots for extrema makes no sense ( at least to me)> @Zepdrix, The second derivative is for convavity

zepdrix (zepdrix):

yes, im aware :) what does the category of "extreme points" include though? I guess I don't recall.

OpenStudy (trojanpoem):

First derivative as the slope is zero in maxima , minima (extrema)

OpenStudy (anonymous):

What would i put for my response for this question?

OpenStudy (trojanpoem):

@Michele_Laino

OpenStudy (anonymous):

Can you help with this one? Same prompt Y = 2/sqrt((1-5)^2)

OpenStudy (trojanpoem):

y = 2/sqrt((1-5)^2) ? y = 2/sqrt(16) = 2/4 = 0.5

OpenStudy (anonymous):

Sorry messed up Y = 2/sqrt(1-5x^2)

OpenStudy (trojanpoem):

What are you trying to do ?

OpenStudy (trojanpoem):

extrema ?

OpenStudy (anonymous):

Same as the first problem

OpenStudy (trojanpoem):

Y' = 2 * -0.5 (1-5x^2)^(-3/2) * -10x = 10x (1-5x^2)^(-3/2) x = 0 1-5x^2 = 0 -5x^2 = -1 x^2 = 1/5 x = +- sqrt(5)/5  Do the rest.

OpenStudy (anonymous):

How did you find this: Y = 2/sqrt((1-5)^2)

OpenStudy (trojanpoem):

I found the derivative using the power rule. You may also use 1/2 sqrt(1-5x^2) * the derivative of what's inside the root.

OpenStudy (anonymous):

Why does the numerator move to the front?

OpenStudy (trojanpoem):

?

OpenStudy (anonymous):

Y = 2/sqrt(1 - 5x^2) Y' = 2(-0.5)....... Why is the 2 from the numerator being multiplied

OpenStudy (trojanpoem):

sqrt = 0.5 1/sqrt = -0.5

OpenStudy (anonymous):

Could have used the quotient rule?

OpenStudy (trojanpoem):

Yeah,

OpenStudy (trojanpoem):

But using the power rule here will help you avoid some unnecessary simplification later.

OpenStudy (michele_laino):

first question: we have no real solutions, so there are not critical points, and since the first derivative is always positive, we can conclude that the function: \(y=x^3-3x^2+4x-4\) is always an increasing function

OpenStudy (anonymous):

Thank you @Michele_Laino @TrojanPoem, after i get 10x/(1-5x^2)^3/2 what do i do from there?

OpenStudy (trojanpoem):

There is critical point when either the derivative is undefined or = zero

OpenStudy (trojanpoem):

so 10x = 0 , 1-5x^2 = 0

OpenStudy (anonymous):

Oh, thank you

OpenStudy (anonymous):

I got sqrt(1/5)

OpenStudy (trojanpoem):

-+ ( when taking roots) sqrt(1)/sqrt(5) = sqrt(5)/5 (convention) no roots in den

OpenStudy (anonymous):

Why cant i square root it as one fraction?

OpenStudy (anonymous):

I get -1 for both positive and negative values of sqrt(5)/5 2/1 - 1 = 0 (Sqrt(5)/5, 0) is my extreme value?

OpenStudy (trojanpoem):

I'ts not one to one function like this y = x^2  It won't be a problem until you're finding inverse (-sqrt(5)/5, -1) , (sqrt(5)/5, -1) ( Ididn't calculate it took your results)